-8x + 14 = -2(4x - 7)
Answer by Theo(13342)

start with:

-8x + 14 = -2 * (4x - 7)

simplify to get:

-8x+ 14 = -8x + 14

the expressions on each side of the equation are identical.

subtract the right side of the equation from both sides of the equation to get:

0 = 0

the variables have disappeared and the equation is true.

that means any value of the variable will also make the equation true.

in other words, there are an infinite number of solutions to this equation.

examples:

when x = 1, 8x + 14 = 22 on both sides of the equation, making them equal to each other.

when x = 100, 8x + 14 = 814 on both sides of the equation, making them equal to each other.